A study is given to college students to determine the semester grades. The results are given in the frequency distribution. Determine the probabilities for each of the following. Use proper symbolism.

Grade

Frequency

A-    To   A+

6

B-    To  B+

12

C-    To  C+

3

D-    To D+

1

          F or WF

2

24

a.    earned better than a D.

b.    A student earned an A or B range grade.

c.     A student earned neither an A nor F range grade.

d.    A student earned at least a C range grade.
Answer #1

Answer)

Total = 24

Probability is given by favorable/total

A)

Favorable = better than D = 3+12+6 = 21

Probability = 21/24 = 0.875

B)

Favorable = A or B = 6+12 = 18

Probability = 18/24 = 0.75

C)

Favorable = neither A nor F = 1+3+12 = 16

Probability = 16/24 = 0.6667

D)

Favorable = at least 6 = 3+12+6 = 21

Probability = 21/24 = 0.875